OpenStudy (anonymous):

greatest common factor for 27,54,and81

OpenStudy (anonymous):

27

OpenStudy (diyadiya):

Factors of 27 = 3 *3 * 3 factors of 54 = 2 * 3* 3 * 3 factors of 81 = 3 *3 * 3 * 3 the common factors in 27,54,81 are 3 * 3 * 3 = 27 so gcf =27

OpenStudy (anonymous):

27